# Machine learning surrogate models of many-body dispersion interactions in polymer melts

### v1

@@ 3.2.1 Effects of extra connections
Regarding the extra connections, we set $p=2$ and $N_{\text{extra}}=50$ for all cases. This introduces only $10\%$ more connections to be encoded, requiring no additional parameters and having a negligible impact on computational cost. However, as shown in Tab. 2, this results in an error decrease of over $20\%$. The enhancement aligns with the many-body nature of the fitting target. By incorporating extra connections closer to the center, even if they are not directly linked to the center atom, the model can more effectively capture finer details of the local atomic environment through enriched feature embeddings of nearby atoms. This is analogous to the coupling of dipole-dipole interactions in MBD via diagonalization. In our neural network model, instead, interactions are “coupled” through the non-linearity of the proposed shifted softplus activation function, as shown in Eq. (14). Nevertheless, since the current results are already highly accurate, further adding more connections yields only marginal improvements, while the computational overhead may become noticeable. The setting chosen in this work adequately meets our need.
Table 2: Effects of extra connections on model performance ($\mathcal{E}_{\text{MARE}}\,(\%)$) for different polymers.
$N_{\text{extra}}$ 0 50
Test C H Cl all C H Cl all
PE 8.45 0.65 - 1.10 5.59 0.41 - 0.71
PP 7.48 0.64 - 1.04 6.38 0.49 - 0.83
PVC 4.39 0.80 1.25 1.28 3.42 0.60 0.98 1.04
@@ 3.2.2 Performance of trainable rbfs
For the rbf encoding, the original SchNet uses a single constant value for all $\gamma_{k}$ and places $\mu_{k}$ at uniform intervals of $0.1\,\text{\AA}$ across the range of interest. This approach is reasonable for general molecular datasets, where structural configurations or densities vary significantly. However, for a polymer cluster, as shown in Fig. 1(c), evenly and finely distributed encoding bases are excessive, given the more consistent spherical geometry, the center-focused target, and the decaying behavior of vdW forces. In this work, we set $N_{\text{rbf}}=100$, sufficient to cover all cases, and make both $\gamma_{k}$ and $\mu_{k}$ trainable to optimize the encoding strategy. While trainable rbfs may not significantly improve model accuracy, since finer fixed bases can adequately encode structure if computational resources are not constrained, they accelerate training convergence and enable the development of extremely lightweight models. These models achieve reasonable accuracy with very few rbfs, as demonstrated in Tab. 3, whereas models with fixed rbfs ($\gamma_{k}=10$ and $\mu_{k}$ uniformly spaced from 00 to $15\,\text{\AA}$) experience a substantial loss in predictive power.
@@
Figure 7: Depiction of the trained rbfs for PE with $N_{\text{rbf}}=100$. The black vertical lines indicate the positions of the rbf centers, $\mu_{k}$, while the red lines represent the averaged interatomic distances for each neighbor relative to the center atom. The blue curves show the evolution of the $\gamma_{k}$ values as a function of distance.
Moreover, with sufficient trainable rbfs, SchNet encodes atomic structures in a more physically meaningful way, enhancing the model’s robustness. As shown in Fig. 7, rbf centers cluster naturally in the short- and mid-range distances, with only a few extending to the long range, where they are weighted by decaying $\gamma_{k}$ values over distance. In the proposed Gaussian-type rbf, the coefficient $\gamma$ controls the width of the function, and a small $\gamma$ allows the rbf to attend a wider range of neighbors. The observed distribution pattern reveals that the model autonomously identifies regions with the most predictive information and averages long-range contributions without requiring fine details. Moreover, at short distances, where nearest-neighbor interactions dominate, an intriguing pattern can be observed: $\mu_{k}$ clusters between interatomic distances to better capture local structural features while avoiding zero gradients, as governed by Eq. (11).
@@ 3.3 Performance for the advanced MBD model
The MBD method introduced in Section 2.1 is commonly referred to as the “plain” MBD. Building on the same methodology, several MBD variants have been developed, including MBD@rsSCS [70], uMBD [71], and MBD-NL [72]. Among these, the MBD@rsSCS version (range-separation self-consistent screening) aligns with the scope of this work that is mainly focused on organic systems. MBD@rsSCS is an advanced formulation designed to seamlessly integrate with DFT-based models, providing a more accurate treatment of polarizability screening across different ranges. This is achieved by applying the range separation technique to the dipole-dipole interaction tensor $\boldsymbol{T}$, allowing a proper derivation of short-range screening and long-range correlation.
@@
To achieve a similar level of accuracy as in the plain MBD case, we would need to increase the size of the model. However, this does not indicate a limitation but rather demonstrates the expected trade-off when addressing a more complex problem. Essentially, this test demonstrates the robustness of our trimmed SchNet in handling altered force distributions, highlighting its adaptability. Furthermore, it motivates the exploration of new techniques to accommodate varying levels of target complexity in the future. Within the scope of this work, where we aim to develop a surrogate model for MBD that integrates with general force field models for molecular simulations, the plain MBD fits better than other MBD variants. Hence, the current trimmed SchNet adequately meets our requirements for this application.
@@ C H all C H all
These results demonstrate that the proposed model generalizes well across polymer chain lengths and minor density variations. The consistent accuracy, even across a wide temperature range, further suggests that the representation of crystallinity in the mixed-temperature dataset is limited, leading to predominantly amorphous structures. This, in turn, may enhance the regularity of the cutoff cluster, thereby simplifying the surrogate modeling task. In future work, we will further investigate the ability of the model to handle pronounced variations in structural regularity and assess its performance in the presence of crystallinity effects.
@@ 3.5 Analysis of the Hessian
Differing from generic deep neural network models, such as fully connected neural networks, SchNet incorporates a more physically meaningful architecture by applying convolutional operations to pairwise atomic connections. This is analogous to the analytical MBD force formulation, where the dipole matrix $\boldsymbol{C}^{\text{MBD}}$ undergoes differentiation and scaling through eigen-decomposition products. Similarly, SchNet effectively scales the convolved atomic connections through activations and computes forces via a gradient operation. Thanks to the infinite continuity of softplus activation functions and TensorFlow’s AD function, it is possible to extract the second derivative of a trained SchNet model, corresponding to the Hessian $\boldsymbol{H}_{1j}$. This tensor serves as a valuable tool for analyzing atomic interactions, as it quantifies the force response on a center atom due to infinitesimal displacements of neighboring atoms within the cutoff range [73]. In Fig. 9, we evaluate the Frobenius norm of $\boldsymbol{H}_{1j}$, referred to as the “condensed Hessian” ${H}_{1j}^{\text{cond}}=\|\boldsymbol{H}_{1j}\|_{\text{F}}$, and compare the condensed Hessian profiles obtained from the analytical MBD model and the trained trimmed SchNet model for a randomly chosen carbon atom in different polymer melts. For all cases, wavy behaviors are observed at long range, attributed to many-body correlation effects and varying atom types over distance. In the analytical Hessian profiles, we find a polynomial decay power below 7, which is lower than the decay power of 8 for the PW case. This further underscores the non-trivial modifications introduced by MBD to the power laws that govern the vdW dispersion interactions [60], while also justifying the need for an extended cutoff distance for MBD force calculations.
@@ 3.6 Towards implementation for MD simulations
We have open-sourced the polymer melt dataset [74] and the TensorFlow code of the trimmed SchNet architecture [59] for users to investigate the surrogate model. However, efficiently integrating MLFFs in MD simulations can be challenging. To address this, we propose using JAX_MD as the MD simulation framework and introducing the trimmed SchNet model as a custom potential implemented with Flax_NNX [75]. Both libraries provide JAX-based APIs [76], enabling a smooth and straightforward integration of the neural network into a standard MD simulation. As demonstrated in the demo provided in our repository, the user simply needs to define a force function that calls the SchNet model to calculate the MBD forces for a given atomic system $\boldsymbol{R}$. Due to VRAM limitations, inference is performed iteratively in a for-loop, processing subsets of the full system. In an NVT simulation of a 9k-atom system on a 32GB Nvidia V100 GPU (with an inference batch size of 3000), the model achieves a speed of 0.02 ms/atom/step, leveraging JAX’s high-performance computing capabilities, whereas a single analytical MBD calculation for one atom in a polymer melt requires approximately 1 s.
As discussed in Section 2.1, the proposed MBD surrogate model serves as attractive components of vdW interactions in the conventional FF setting. Therefore, in practice, one must determine an appropriate repulsive term to complete the vdW model for MD simulations. However, this may require reparameterization of the existing FF. For example, the TraPPE FF [77], which is commonly used for polymers built from alkane chains, models the vdW interaction with 12-6 LJ potential and employs shifted interaction sites for hydrogen atoms. In such a case, directly substituting the attractive part with the trained SchNet model is not feasible. A careful compatibility assessment and reparameterization of the repulsive part are necessary for effective coupling. Developing a general solution for this coupling challenge will be part of our future work.
@@ 4 Conclusions and future work
In this paper, we introduced a trimmed SchNet architecture as a surrogate model specifically designed for predicting MBD forces in polymer melts. Despite the widely recognized success of MBD over PW methods in accurately modeling many-body vdW effects, and the rapid growth of MLFFs, our work is the first explicit attempt to develop an ML surrogate for MBD tailored for large-scale molecular simulations. Polymer melts were chosen as the initial application due to their strong dependence on vdW interactions, and their structural characteristic guided both our modeling strategy and architecture design.
@@
With the proposed trimmed SchNet model, one can already integrate it into MD simulations of large-scale polymer melts, provided that it is properly coupled with an existing FF model for other interactions in the system. Our immediate future work will focus on this integration, with particular interest in potential emergent MBD effects that could lead to significant differences in polymer melt morphology and mechanical properties compared to the results from PW vdW models.

### v2

@@ 3.2.1 Effects of extra connections
Regarding the extra connections, we use $p=2$ neareast neighbors and $N_{\text{extra}}=50$ for all cases. This introduces only $10\%$ more connections to be encoded, requiring no additional parameters and having a negligible impact on computational cost. As shown in Tab. 2, $p=2$ provides the most consistent accuracy across all polymer types, offering notable improvements over no extra connection ($p=0$) or only $p=1$ nearest neighbor. While $p=3$ may yield slight gains for certain cases, the changes are marginal under the same training strategy. The enhancement from these extra connections aligns with the many-body nature of the fitting target: by incorporating extra connections closer to the center, even if not directly linked to it, the model can more effectively capture finer details of the local atomic environment through enriched feature embeddings of nearby atoms. This is analogous to the coupling of dipole–dipole interactions in MBD via diagonalization. In our neural network model, instead, interactions are “coupled” through the non-linearity of the proposed shifted softplus activation function, as shown in Eq. (14).
Table 2: Effects of extra connections on model performance ($\mathcal{E}_{\text{MARE}}\,(\%)$) for different polymers. Results are shown for varying $p$ values (number of nearest neighbors) with $N_{\text{extra}}$ fixed at 50.
p 0 1 2 3
Test C H Cl all C H Cl all C H Cl all C H Cl all
PE 8.45 0.65 - 1.10 6.67 0.46 - 0.81 5.59 0.41 - 0.71 5.43 0.46 - 0.74
$P$ 7.48 0.64 - 1.04 10.22 0.54 - 1.10 6.38 0.49 - 0.83 6.20 0.56 - 0.89
PVC 4.39 0.80 1.25 1.28 3.82 0.55 0.93 0.97 3.42 0.60 0.98 1.04 3.34 0.57 0.99 0.95
@@ 3.2.2 Performance of trainable rbfs
For the rbf encoding, the original SchNet uses a single constant value for all $\gamma_{k}$ and places $\mu_{k}$ at uniform intervals of $0.1\,\text{\AA }$ across the range of interest. This approach is reasonable for general molecular datasets, where structural configurations or densities vary significantly. However, for a polymer cluster, as shown in Fig. 1(c), evenly and finely distributed encoding bases are excessive, given the more consistent spherical geometry, the center-focused target, and the decaying behavior of vdW forces. In this work, we set $N_{\text{rbf}}=100$, sufficient to cover all cases, and make both $\gamma_{k}$ and $\mu_{k}$ trainable to optimize the encoding strategy. While trainable rbfs may not significantly improve model accuracy for large $N_{\text{rbf}}$, since finer fixed bases can adequately encode structure if computational resources are not constrained, they can accelerate training convergence for smaller $N_{\text{rbf}}$. In particular, for $N_{\text{rbf}}=10$ and $20$, the trainable rbfs exhibit a much faster loss decay and achieve lower final training loss compared to fixed rbs, as shown in Fig. 7. Here, the fixed rbf uses $\gamma_{k}=10$ and $\mu_{k}$ uniformly spaced between 0 and $15\,\text{\AA }$, which also serves as the initialization for the trainable case. This ability to maintain reasonable accuracy with very few rbfs enables the development of extremely lightweight models, as demonstrated in Tab. 3, whereas models with fixed rbfs under the same settings experience a substantial loss in predictive power.
Figure 7: Training convergence of the model on PE dataset using different rbf encoding strategies. Three $N_{\text{rbf}}$ values are considered. Solid curves represent trainable rbfs and dashed curves represent fixed rbfs.
@@
Moreover, with sufficient trainable rbfs, SchNet encodes atomic structures in a more physically meaningful way, enhancing the model’s robustness. As shown in Fig. 8, rbf centers cluster naturally in the short- and mid-range distances (see black vertical lines), with only a few extending to the long range, where they are weighted by decaying $\gamma_{k}$ values over distance (see the blue line). In the proposed Gaussian-type rbf, the coefficient $\gamma$ controls the width of the function, and a small $\gamma$ allows the rbf to attend a wider range of neighbors. The observed distribution pattern reveals that the model autonomously identifies regions with the most predictive information and averages long-range contributions without requiring fine details. Notably, at short distances, where nearest-neighbor interactions dominate, an intriguing pattern can be observed: rbf centers cluster between interatomic distances (red vertical lines) to better capture local structural features while avoiding zero gradients, as governed by Eq. (11).
Figure 8: Visualization of the trained rbfs for PE with $N_{\text{rbf}}=100$. The x-axis shows distance relative to the center atom, and the y-axis shows rbf coefficient $\gamma_{k}$. Black vertical lines indicate the learned positions of the rbf centers $\mu_{k}$, red lines mark the averaged interatomic distances for each neighbor relative to the center atom, and the blue curve illustrate the corresponding $\gamma_{k}$ values, controlling the width of each basis function.
@@ 3.3 Performance for the advanced MBD model
The MBD method introduced in Section 2.1 is commonly referred to as the “plain” MBD. Building on the same methodology, several MBD variants have been developed, including MBD@rsSCS [3], universal MBD (uMBD) [47], and nonlocal MBD (MBD-NL) [38]. Among these, the MBD@rsSCS version (range-separation self-consistent screening) aligns with the scope of this work that is mainly focused on organic systems. MBD@rsSCS is an advanced formulation designed to seamlessly integrate with Density Functional Theory (DFT)-based models, providing a more accurate treatment of polarizability screening across different ranges. This is achieved by applying the range separation technique to the dipole-dipole interaction tensor $\boldsymbol{T}$, allowing a proper derivation of short-range screening and long-range correlation.
@@
The reduced performance of the model on MBD@rsSCS can be attributed to the increased intrinsic complexity of this variant. Compared to the plain MBD, MBD@rsSCS introduces short-range self-consistent screening and modifies the dipole interaction tensor at long range [3], both of which significantly alter the overall force landscape and force magnitudes. These effects likely make the learning task more challenging. To achieve a similar level of accuracy as in the plain MBD case, we would need to increase the size of the model or apply dedicated hyperparameter tuning to better capture the more complex force distribution. However, this does not indicate a limitation but rather demonstrates the expected trade-off when addressing a more complex problem. Essentially, this test demonstrates the robustness of our trimmed SchNet in handling altered force distributions, highlighting its adaptability. Furthermore, it motivates the exploration of new techniques to accommodate varying levels of target complexity in the future. Within the scope of this work, where we aim to develop a surrogate model for MBD that integrates with general force field models for molecular simulations, the plain MBD fits better than other MBD variants. Hence, the current trimmed SchNet adequately meets our requirements for this application.
@@ C H all C H all
These results demonstrate that the proposed model generalizes well across polymer chain lengths and minor density variations. The consistent accuracy, even across a wide temperature range, further suggests that the representation of crystallinity in the mixed-temperature dataset is limited, leading to predominantly amorphous structures. This, in turn, may enhance the regularity of the cutoff cluster, thereby simplifying the surrogate modeling task. We acknowledge that a more rigorous validation would require testing on systems with pronounced crystallinity, which could be achieved through longer equilibration or dedicated crystalline sampling. Such systems will be included in future work to evaluate the model’s robustness under reduced structural regularity and the presence of ordered domains.
@@ 3.5 Analysis of the Hessian
Differing from generic deep neural network models, such as fully connected neural networks, SchNet incorporates a more physically meaningful architecture by applying convolutional operations to pairwise atomic connections. This is analogous to the analytical MBD force formulation, where the dipole matrix $\boldsymbol{C}^{\text{MBD}}$ undergoes differentiation and scaling through eigen-decomposition products. Similarly, SchNet effectively scales the convolved atomic connections through activations and computes forces via a gradient operation. Thanks to the infinite continuity of softplus activation functions and TensorFlow’s AD function, it is possible to extract the second derivative of a trained SchNet model, corresponding to the Hessian $\boldsymbol{H}_{1j}$. This tensor serves as a valuable tool for analyzing atomic interactions, as it quantifies the force response on a center atom due to infinitesimal displacements of neighboring atoms within the cutoff range [35]. In Fig. 10, we evaluate the Frobenius norm of $\boldsymbol{H}_{1j}$, referred to as the “condensed Hessian” ${H}_{1j}^{\text{cond}}=\|\boldsymbol{H}_{1j}\|_{\text{F}}$, and compare the condensed Hessian profiles obtained from the analytical MBD model and the trained trimmed SchNet model for a randomly chosen carbon atom in different polymer melts. This condensed form serves as a scalar representation for the effective force constants associated with MBD interactions, offering a physically intuitive and easily visualizable measure of how MBD stiffness decays with interatomic distance. For all cases, wavy behaviors are observed at long range, attributed to many-body correlation effects and varying atom types over distance. In the analytical Hessian profiles, we find a polynomial decay power below 7, which is lower than the decay power of 8 for the PW case. This further underscores the non-trivial modifications introduced by MBD to the power laws that govern the vdW dispersion interactions [2], while also justifying the need for an extended cutoff distance for MBD force calculations.
@@ 3.6 Towards implementation for MD simulations
We have open-sourced the polymer melt dataset [72] and the TensorFlow code of the trimmed SchNet architecture [71] to facilitate future research on surrogate modeling of MBD interactions. One of the most direct use of our surrogate model would be its integration in MD simulations. However, such an integration can be challenging, and in this section we discuss possible approaches to do it efficiently.
@@
Technically, as the MD simulation framework we propose using JAX_MD, with the trimmed SchNet model introduced as a custom potential implemented with Flax_NNX [37]. Both libraries provide JAX-based APIs [13], enabling a smooth and straightforward integration of the neural network into a standard MD simulation. As demonstrated in the demo provided in our repository, the user simply needs to define a force function that calls the SchNet model to calculate the MBD forces for a given atomic system $\boldsymbol{R}$. Due to VRAM limitations, inference is performed iteratively in a for-loop, processing subsets of the full system.
To estimate the computational performance of the model, we have carried out an NVT simulation of a 9k-atom system on a 32GB Nvidia V100 GPU (with an inference batch size of 3000). The model predictions achieve a speed of 0.02 ms/atom/step, leveraging JAX’s high-performance computing capabilities. This is several order of magnitude faster to a single analytical MBD calculation for one atom in a polymer melt, which requires approximately 1 s.
Figure 11: Time evolution of the radius of gyration $R_{g}(t)$ for six polymer chains in a PE melt during an NVT MD simulation using the trimmed SchNet MBD surrogate. Each curve corresponds to one polymer chain and represents an average over multiple independent random-seed simulations.
To demonstrate that the proposed surrogate can be stably integrated into an MD workflow, we performed a preliminary MD test using a practical FF coupling strategy for polymer melts. As discussed in Section 2.1, the MBD surrogate model serves as the attractive component of the vdW interaction, and must therefore be combined with additional interaction terms to enable MD simulations. In this context, the surrogate is coupled to TraPPE [14], a classical FF model for polymers, following the implementation strategy detailed in [73]. The simulation is conducted at 300 K in the NVT ensemble for a PE system with approximately 9k atoms, consisting six polymer chains. As a simple and widely used structural observable, the time evolution of the radius of gyration $R_{g}(t)$ is monitored for each chain for around 150 picoseconds. The resulting trajectories shown in Fig. 11 exhibit smooth and bounded behavior over a relatively long simulated time window, indicating that the surrogate model behaves stably when used in a standard MD workflow. The purpose of this test is solely to demonstrate numerical stability and practical integrability of the model, rather than to validate polymer structural or thermodynamic properties.
The above mentioned practical coupling is in general not sufficient. Deploying the MBD surrogate as part of a complete and quantitatively predictive FF for polymer melts requires the inclusion of bonded interactions and a consistent short-range repulsive component for the vdW interaction. In principle, the surrogate can be coupled to an existing or lightly tuned LJ repulsive term within a classical FF to enable rapid, qualitative investigations of MBD effects relative to the PW model. For rigorous, quantitative coupling, however, the repulsive parameters should be refitted using the same physical targets employed in the original force field development. For example, in TraPPE, LJ parameters are calibrated to reproduce vapor–liquid phase equilibria, and replacing the attractive component with the MBD surrogate would naturally be accompanied by re-calibrating the repulsive term to preserve these target properties.
In even longer perspective, a more physically consistent integration can be achieved by coupling the surrogate to a short-range MLFF such as SO3LR or GEMS. These general-purpose MLFFs already model bonded and short-range non-bonded interactions at high fidelity by fitting to DFT-level data, whereas explicit long-range MBD effects are not yet incorporated in these approaches. Moreover, in the context of the range-separation concept as discussed in [70], DFT-based models inherently lack a description of long-range correlation and require a vdW dispersion complement. This makes coupling to the MBD surrogate both smooth and physically consistent. Such a combination would yield a complete MLFF with accurate short-range physics and long-range MBD interactions. Developing and validating this coupling, including tests on structural, dynamic, and thermodynamic properties from MD simulations, are in progress but lie beyond the scope of the present manuscript.
@@ 4 Conclusions and future work
In this paper, we introduced a trimmed SchNet architecture as a surrogate model specifically designed for predicting MBD forces in polymer melts. Despite the widely recognized success of MBD over PW methods in accurately modeling many-body vdW effects, and the rapid growth of MLFFs, dedicated ML surrogates for MBD remain largely unexplored. Our work addresses this gap by presenting a model tailored for MBD in large-scale molecular simulations. Polymer melts were chosen as the initial application due to their strong dependence on vdW interactions, and their structural characteristic guided both our modeling strategy and architecture design.
@@
With the proposed trimmed SchNet model, one can already integrate it into MD simulations of large-scale polymer melts, provided that it is consistently coupled with an appropriate FF model for the remaining interactions in the system. While the present work demonstrates the numerical stability and practical feasibility of such integration, achieving a rigorous and quantitatively predictive coupling requires further effort. Our immediate future work will focus on this direction, with particular interest in potential emergent MBD effects that could lead to significant differences in polymer melt morphology and mechanical properties compared to the results from PW vdW models.